OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

show answer step by step please :Dplease help!!!

OpenStudy (lgbasallote):

use the formula \[\large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\] need any more help?

OpenStudy (anonymous):

yup :/

OpenStudy (lgbasallote):

okay...y1 means y coordinate of the first ordered pair x1 is the x coordinate of the first ordered pair x2 is the x coordinate of the second ordered pair y2 is the y coordinate of the second ordered pair now do you know how to do it?

OpenStudy (anonymous):

is it -1/3?

OpenStudy (lgbasallote):

denominator would be 5 - (-2) = 5 + 2

OpenStudy (lgbasallote):

your numerator is correct

OpenStudy (anonymous):

so...-1/7

OpenStudy (anonymous):

what do i do next?.. dont i have to put it x+yb=b form?

OpenStudy (lgbasallote):

-1/7 is just the slope the y2 - y1/x2 - x1 part remember i wrote y - y1 and x - x1

OpenStudy (zepp):

@seashell, sorry, I was helping somebody else when you mentioned me :s

OpenStudy (anonymous):

its ok:D

OpenStudy (lgbasallote):

\[y - y_1 = -\frac{1}{7} (x - x_1)\] what's your y1 and x!?

OpenStudy (anonymous):

2 and -2

OpenStudy (lgbasallote):

2 is y1 -2 is x1 \[y - 2 = -\frac{1}{7} (x +2 )\] ^x+ 2 because x - (-2) = x + 2 now multiply both sides by 7 to get rid of the fraction

OpenStudy (zepp):

http://openstudy.com/users/zepp#/updates/4fa3474fe4b029e9dc34125e Refer to this guide! :D

OpenStudy (zepp):

Situation 4 :)

OpenStudy (anonymous):

how do i do that...

OpenStudy (lgbasallote):

if you multiply both sides by 7 you'll get \[7(y - 2) = -1(x +2)\]

OpenStudy (lgbasallote):

simplify that by distribution